Is this the proper way to word these two problems ???
3-(2+3x)
Answer Choices:
A: 1-3x
B: 1+3x<--- This was my son and his tutor's choice
C: 5+3x

8d-(4d+2)
Answer Choices:
A:4d+2<--- This was my son and his tutor's choice.
B:4d-2
C:2d
My son is working on an Internet course from BYU (Brigham Young University). The instructor says that my sons choices were wrong. I believe that BYU is wrong, can you please clarify BYU's reasoning???

Signed, concerned Dad

Answer by venugopalramana(3286)

SEE MY COMMENTS BELOW..
Is this the proper way to word these two problems ???
3-(2+3x)..OPEN THE BRACKETS.THE MINUS OR -VE SIGN INFRONT OF BRACKET APPLIES TO ALL TERMS WITHIN THE BRACKET.INFACT WE HAVE TO MULTIPLY EACH AND EVERY TERM WITHIN THE BRACKET WITH THE COEFFICIENT OUT SIDE THE BRACKET WHICH IS -1 HERE..SO WE GET
3-1*2-1*3X=3-2-3X=1-3X.....IF YOU ARE STILL IN DOUBT ,BEST WAY IS TO IMAGINE SOME VALUE FOR X APPLY AND SEE...SAY X=1...WE GET....3-(2+3*1)=3-(2+3)=3-(5)
=-2......1-3X ALSO GIVES SAME ANSWER =1-3*1=1-3=-2...WHERE AS 1+3X GIVES 1+3*1=1+3=4 WHICH IS NOT CORRECT.
Answer Choices:
A: 1-3x
B: 1+3x<--- This was my son and his tutor's choice
C: 5+3x
8d-(4d+2)..SAME LOGIC AS ABOVE APPLIES HERE.WE GET 8D-4D-2=4D-2..YOU CAN CHECH BY PUTTING D=1
Answer Choices:
A:4d+2<--- This was my son and his tutor's choice.
B:4d-2
C:2d
My son is working on an Internet course from BYU (Brigham Young University). The instructor says that my sons choices were wrong. I believe that BYU is wrong, can you please clarify BYU's reasoning???
Signed, concerned Dad